OpenStudy (anonymous):

joseph who is paid time and half for hours worked over 40 had gross weekly wages of $442 for working 48 hrs. what is his regular hourly pay?

OpenStudy (anonymous):

Let x = normal rate. Since he makes time and a half for ever hour over 40, his overtime rate is 1.5(x). So if he worked 48 hours, he was paid 40 hours at x dollars an hour, and 8 hours at 1.5x dollars an hour. He worked a total of 48 hours, so his total pay was 40(x) + 8(1.5x) = 442. Multiplying out, we get 52x = 442, or x equals 8.5. So his regular hourly pay is 8.50.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

You're very welcome!

OpenStudy (anonymous):

i have another question, find linear formula if 2 box cost $4 and 4 box cost $6

OpenStudy (anonymous):

I'm not sure what you're asking. If you are asking to make them into a line, then you could do it as follows. Since we know 2 boxes cost 4 dollars, and 4 boxes cost 6 dollars, then to make this a linear formula, we must make it in the form y = mx + b. We know two of the points, namely (2, 2) and (4, 6) since the boxes and their costs are given. I've assigned x to the boxes, and y to the cost. Plugging in both points in y = mx + b, we have 2 = 2m + b, and 6 = 4m + b. m is the slope, and is equal to the change in y over the change in x. Change in Y = 4 when x changes by 2 (X goes up 2 boxes, Y goes up 4 dollars), so the slope is positive 2. So we know have 2 = 4 + b and 6 = 8 + b, getting b = -2. Since m = 2, and b = -2, y = mx + b ==> y = 2x - 2.